OpenStudy (anonymous):

2x+y=10 leave y by itself

OpenStudy (anonymous):

subtract 2x from each side

OpenStudy (anonymous):

y = -2 (x - 5)

OpenStudy (anonymous):

@pleasehelpasap y = 2(x - 5) -.-

OpenStudy (anonymous):

typo i meant y = 2(x -5) @gaara438125 leave me alone and stop stalking me!

OpenStudy (anonymous):

@pleasehelpasap stop breaking the rules...